solve
0.2a - 0.9a + 7.5 = 10.3

Answer by checkley77(12844)

0.2a - 0.9a + 7.5 = 10.3
-.7a=10.3-7.5
-.7a=2.8
a=2.8/-.7
a=-4 answer.
proof
.2*-4-.9*-.4+7.5=10.3
-.8+3.6+7.5=10.3
10.3=10.3